OpenStudy (anonymous):

if f(x)=-x^2-3 then find f(a+1)

OpenStudy (anonymous):

just substitute a+1 instead of x and simplify

OpenStudy (anonymous):

f(a+1)=-(a+1)^2 - 3=-(a^2 + 2a + 1) - 3= -a^2 - 2a - 1 - 1= -a^2 - 2a - 4 so f(a+1)=-a^2-2a-4

OpenStudy (anonymous):

when you put -a^2-2a-1-1 did you mean -a^2-2a-1-3? i dont understand where that second -1 came from

OpenStudy (anonymous):

oh yes. sorry. i meant -1-3

OpenStudy (anonymous):

Okay! thank you for the help!